# This version: August 29, 2025 )

### v3

@@ Abstract
A growing number of central authorities use assignment mechanisms to allocate students to schools in a way that reflects student preferences and school priorities. However, most real-world mechanisms incentivize students to strategically misreport their preferences. Misreporting complicates the identification of causal parameters that depend on true preferences, which are necessary inputs for a broad class of counterfactual analyses. In this paper, we provide an identification approach that is robust to strategic misreporting and derive sharp bounds on causal effects of school assignment on future outcomes. Our approach applies to any mechanism as long as there exist placement scores and cutoffs that characterize that mechanism’s allocation rule. We use data from a deferred acceptance mechanism that assigns students to more than 1,000 university–major combinations in Chile. Matching theory predicts and empirical evidence suggests that students behave strategically in Chile because they face constraints on their submission of preferences and have good a priori information on the schools they will have access to. Our bounds are informative enough to reveal significant heterogeneity in graduation success with respect to preferences and school assignment.
@@ 1 Introduction
One of the most important decisions students make is their choice of field and institution of education. Identification of the impact of such choices on future outcomes is a critical step in the study of this decision process and of public education policies. The causal effects of schooling may vary widely across economic agents because of heterogeneous skills and preferences. Moreover, individuals’ expectations about potential returns may prompt them to choose schools strategically. Heterogeneity and selection make identification of causal effects very challenging, especially when students face a large number of unordered options. On the positive side, a growing number of schools use centralized assignment mechanisms, which produce credible instruments based on the discontinuities generated by the assignment of comparable students to different schools (Kirkeboen et al. (2016) and Abdulkadiroglu et al. (2022)).
@@
Many centralized school assignment mechanisms effectively amount to a quasi-experimental design where two groups of individuals who share similar scores are assigned to different schools based on how their scores relate to admission cutoffs. The assignment in a matching characterized by such cutoffs depends on the student’s preferences over feasible schools. Unlike in a typical regression discontinuity (RD) design, in this setting, students on the same side of the cutoff do not necessarily receive the same assignment. For example, individuals with similar scores just above a certain cutoff may all prefer to go to the same “school jj” but could have very different second-best options if they fall on the other side of the cutoff. In such a context, Kirkeboen et al. (2016) construct comparable groups of individuals near a cutoff by conditioning on local preferences—that is, by selecting students whose preferences yield identical first- and second-best options if they fall, respectively, above and below that cutoff. Controlling for local preferences that equal a pair of schools, e.g., $(j,k)$, allows the RD to identify the causal effect of a change in the school assignment from $k$ to $j$, averaged over individuals who prefer $j$ over $k$.
@@
Most real-world school assignment mechanisms create incentives for students to misreport their true preferences. Agarwal and Somaini (2018) and Fack et al. (2019) provide thorough discussions with several real-world examples of this. The empirical literature on school choice typically engages in counterfactual policy evaluations and an important class of such counterfactuals requires identification of causal parameters that are conditional on true preferences (Artemov et al. (2023), Section V.B). The identification of such causal parameters is the primary goal of our paper and we propose a novel RD strategy that controls for true local preferences. The challenge resides on the fact that true preferences are unobserved but may be partially identified under general forms of misreporting behavior.
@@
This paper derives sharp bounds for causal effects of school assignment on future outcomes in mechanisms with cutoff characterization and strategic student behavior. We devise a two-step identification approach that is robust to strategic reporting of preferences. In the first step, the researcher partially identifies local preferences and constructs local preference sets for each student. We provide several tools for constructing these sets in the context of a student-proposing deferred acceptance (DA) mechanism where constraints on the submitted preferences lead to strategic behavior. Outside such a context, researchers may employ alternative tools to partially identify local preferences, as the second step of our procedure does not require a particular method to be used in the first step. For example, the identification method of Agarwal and Somaini (2018) applies to a general class of mechanisms with cutoff representation that includes variants of the DA mechanism, Boston, First Preferences First, Chinese Parallel, etc. Finally, in the second step, the researcher employs an RD identification strategy that controls for local preference sets and partially identifies the causal effects of school assignment.
@@
Strategic behavior among students depends on the characteristics of the assignment mechanism. A mechanism is said to be strategy proof if submitting true preferences is a weakly dominant strategy for all students. For example, Dubins and Freedman (1981) demonstrate that the DA mechanism is strategy proof. However, this result breaks down when the mechanism imposes constraints on the preferences that students can submit. In many real-world school assignment mechanisms, the number of schools is too large for students to feasibly rank all schools. The central authorities running these systems may either limit the number of schools that students may rank or impose costs on the basis of the number of schools submitted (see Table 1 Panel B by Fack et al. (2019) for examples).
@@
Our first-step tools for partial identification of local preferences naturally require assumptions on students’ strategic behavior. We motivate our assumptions following the important contributions of Haeringer and Klijn (2009). They study a game where students submit constrained preference rankings and a central mechanism allocates the students to schools. One of their important findings is that it is rational for students to submit partial orders of their true preferences in some mechanisms. Specifically, suppose that a mechanism is strategy proof when students are free to rank any number of schools, as under, e.g., the unconstrained DA or Top Trading Cycles (TTC) mechanisms. Then, if the preference rankings are constrained to having at most $K$ schools, a student can do no better than selecting $K$ schools among her acceptable schools and ranking them according to her true preferences.
@@
The key assumption for our first-step tools is that students submit only partial orders of their preferences. This feature is in addition to the cutoff characterization of the matching, which we assume throughout the paper. Cutoff characterization means that a student is matched to her best feasible school, where “best” is defined according to her true preferences and a school is feasible if the student’s placement score clears that school’s admission cutoff. Our assumption about cutoff characterization is satisfied when the matching outcome is stable (Azevedo and Leshno, 2016). Stability means that each student is matched to an acceptable school and all slots in preferred schools have been filled with people who have better placement scores. In constrained DA and TTC mechanisms, stability occurs in Nash equilibrium of the preference revelation game under appropriate conditions on the placement scores (Theorems 6.3 and 6.4, Haeringer and Klijn (2009)). Stability occurs in Nash equilibrium without restrictions on the scores in the constrained serial dictatorship (SD) mechanism, which is a particular case of DA. We characterize sharp local preference sets for every individual that are compatible with the observed data and these model assumptions. Our random local preference sets contain the true local preference random variable with probability one. We also show a way to shrink these sets by imposing assumptions on students’ expectations regarding the outcome of the match and by assuming students maximize expected utility, in line with Agarwal and Somaini (2018).
@@
The second step of our approach relies on the local preference sets constructed in the first step with either our method or an alternative method. Given interest in a pair of schools $(j,k)$, we select all individuals whose local preference sets contain $(j,k)$ and whose placement scores are close to the cutoff for admission at school $j$. This subpopulation of individuals contains all individuals whose true local preferences equal $(j,k)$, but also other individuals. The average outcome in the subpopulation equals a weighted average of two averages: first, the average outcome for individuals with true local preferences $(j,k)$, which is interesting for the identification of causal effects, and second, the average outcome for individuals with true local preferences that differ from $(j,k)$. We do not know which individuals have preferences $(j,k)$, but we do characterize sharp bounds on the proportion of such individuals in the subpopulation using the random sets constructed in the first step. Thus, our setting fits the identification problem with corrupted data studied by Horowitz and Manski (1995). This method allows us to derive closed-form bounds on the first out of the two average outcomes above, which then leads to bounds on the average causal effects. This closed-form approach offers some intuition on when we can expect the bounds to be informative about or equal to the actual average causal effect (i.e., point identification). Although practical and intuitive, these closed-form bounds may not be sharp. Thus, building on Molinari (2020) and using random set theory, we characterize sharp bounds that are numerically computable when outcomes take finitely many values.
@@
Methods combining RD identification with school matching data have been popular among applied and theoretical researchers in economics (Jackson, 2010; Bertanha, 2020). To the best of our knowledge, our paper is the first to prove RD identification of returns of school assignment in matching mechanisms with strategically reported preferences. We note that our two-step approach differs from the usual control function approach because our first step partially identifies the control variable instead of point-identifying it as in the usual approach. For this reason, we call it control mapping approach. Various other examples in econometrics feature this structure, see Han and Kaido (2025). Our paper unifies and complements two branches of the literature. One branch features the methods proposed by Agarwal and Somaini (2018) and Fack et al. (2019) that take into account strategic reporting and identify students’ true preferences, but they do not focus on causal effects of school matches on future outcomes. The other branch features the methods of Kirkeboen et al. (2016) and Abdulkadiroglu et al. (2022) that identify the causal effects of different assignments but control for reported instead of true preferences. An RD strategy that controls for reported preferences identifies causal parameters that could be useful for a class of counterfactuals that differs from the one we consider in this paper. It is important to note, however, that in some empirical contexts, agents may have precise ex ante information about ex post matching cutoffs at the time when they choose preferences to report. In this case, misreporting behavior may arise near a cutoff and this confounds identification in an RD that controls for reported preferences; we find that controlling for true preferences makes the RD strategy robust to this possibility. In the same branch of the literature, Chen (2023) provides a comprehensive set of causal parameters that are identified when assignment is based both on lottery- and RD-driven variation.
We apply our two-step identification strategy to matching data from Chile. Chile has a centralized DA mechanism that assigns students to university–major pairs. In 2010, 88,000 students ranked at most eight university–major pairs out of a total of 1,092 options available. Thus, the mechanism constrains students’ preference rankings, and students have incentives to behave strategically. The methods proposed by Kirkeboen et al. (2016) and Abdulkadiroglu et al. (2022) are not directly applicable in the Chilean case, even if all students report their preferences truthfully, as we explain below.
The methods of Kirkeboen et al. (2016) apply to the SD mechanism—a particular case of the DA mechanism in which all schools utilize the same placement score. In the SD case, the counterfactual set of schools for students just above or just below a cutoff does not vary across students. For example, for students just above a cutoff, their counterfactual set includes all schools whose cutoffs are lower than the cutoff in question. The same does not apply in DA, the mechanism used in Chile. Students have multiple placement scores, and the set of feasible schools may vary widely across students. Defining the counterfactual set is an important step in the RD identification strategy because local preferences are defined over these sets, which then become the control variable in the RD strategy.
Abdulkadiroglu et al. (2022) propose a solution to the DA counterfactual problem by constructing a propensity score control variable. They study New York City public high schools, where placement scores are functions of integer priority scores plus a continuously distributed variable with full support. This particular structure of school priorities does not correspond to the Chilean case, where program-specific placement scores are computed as functions of five primitive scores: math, language, history, science, and an average score from high school. These functions are different across programs and sometimes nonlinear. This setting requires the counterfactual set of schools to be carefully defined such that controlling for local preferences does not violate the continuity assumptions required by RD. We therefore propose a general method that applies to such empirical contexts and leads to point identification if students are truthful but partial identification otherwise.
@@
We illustrate our methods using data from college applications in Chile, beginning by documenting clear evidence of strategic behavior among applicants. The analysis proceeds in two parts. First, we estimate average structural functions in the context of the Bachillerato de Ingreso Común at the University of Chile—a selective entry program with a large applicant pool that allows students to explore multiple disciplines before declaring a major, and is often viewed as an alternative route into competitive programs. We examine whether assignment to this program affects two outcomes: graduating from one’s original local first-best program $j$ and graduating from a top university. The results reveal substantial heterogeneity across $j$, indicating that students’ preferences matter for graduation outcomes and are correlated with unobserved factors such as effort or ability not captured by test scores. Second, we estimate treatment effects focusing on Medicine at PUC Santiago, a program whose high selectivity, popularity, and predictable cutoff make strategic behavior especially likely. We measure the effect of being assigned to Medicine rather than to common next-best alternatives $k$ on the two graduation outcomes above. In all cases, our preferred bounds identify a positive effect, with magnitudes that vary across alternatives $k$.

### v4

# This version: March 9th, 2026 )

@@ Abstract
A growing number of authorities use mechanisms to allocate students to schools in a way that reflects student preferences and school priorities. However, most real-world mechanisms incentivize students to strategically misreport their preferences. Misreporting complicates the identification of causal parameters that depend on true preferences, which are necessary inputs for a broad class of counterfactual analyses. We provide an identification approach robust to misreporting and derive sharp bounds on causal effects of school assignment. Our approach applies to allocation rules characterized by placement scores and cutoffs. We use data from a deferred acceptance mechanism that assigns students to university programs in Chile. Matching theory predicts and empirical evidence shows that students behave strategically in Chile because they face constraints on preference submission and have good prior information about school accessibility. Our bounds are informative enough to reveal significant heterogeneity in graduation success with respect to preferences and school assignment.
@@ 1 Introduction
One of the most important decisions students make is their choice of field and institution of education. Identification of the impact of such choices on future outcomes is a critical step in the study of this decision process and of public education policies. The causal effects of schooling may vary widely across economic agents because of heterogeneous skills and preferences. Moreover, individuals’ expectations about potential returns may prompt them to choose schools strategically. Heterogeneity and selection make identification of causal effects very challenging, especially when students face a large number of unordered options. On the positive side, a growing number of schools use centralized assignment mechanisms, which produce credible instruments based on the discontinuities generated by the assignment of comparable students to different schools (kirkeboen2016 and abdul_ecta2022).
@@
Many centralized school assignment mechanisms effectively amount to a quasi-experimental design where two groups of individuals who share similar scores are assigned to different schools based on how their scores relate to admission cutoffs. The assignment in a matching characterized by such cutoffs depends on the student’s preferences over feasible schools. Unlike in a typical regression discontinuity (RD) design, in this setting, students on the same side of the cutoff do not necessarily receive the same assignment. For example, individuals with similar scores just above a certain cutoff may all prefer to go to the same “school jj” but could have very different second-best options if they fall on the other side of the cutoff. In such a context, kirkeboen2016 construct comparable groups of individuals near a cutoff by conditioning on local preferences—that is, by selecting students whose preferences yield identical first- and second-best options if they fall, respectively, above and below that cutoff. Controlling for local preferences that equal a pair of schools, e.g., $(j,k)$, allows the RD to identify the causal effect of a change in the school assignment from $k$ to $j$, averaged over individuals who prefer $j$ over $k$.
@@
Most real-world school assignment mechanisms create incentives for students to misreport their true preferences. agarwal2018 and fack2019 provide thorough discussions with several real-world examples of this. Subsequent empirical work have employed their methods to estimate true preferences or test for strategic behavior, for example, campos_impact_2024, andersson_beyond_2026, and many others. The empirical literature on school choice typically engages in counterfactual policy evaluations and an important class of such counterfactuals requires identification of causal parameters that are conditional on true preferences (Artemov2023, Section V.B). The identification of such causal parameters is the primary goal of our paper and we propose a novel RD strategy that controls for true local preferences. The challenge resides on the fact that true preferences are unobserved but may be partially identified under general forms of misreporting behavior.
@@
This paper derives sharp bounds for causal effects of school assignment on future outcomes in mechanisms with cutoff characterization and strategic student behavior. We devise a two-step identification approach that is robust to strategic reporting of preferences. In the first step, the researcher partially identifies local preferences and constructs local preference sets for each student. We provide several tools for constructing these sets in the context of a student-proposing deferred acceptance (DA) mechanism where constraints on the submitted preferences lead to strategic behavior. Outside such a context, researchers may employ alternative tools to partially identify local preferences, as the second step of our procedure does not require a particular method to be used in the first step. For example, the identification method of agarwal2018 applies to a general class of mechanisms with cutoff representation that includes variants of the DA mechanism, Boston, First Preferences First, Chinese Parallel, etc. Finally, in the second step, the researcher employs an RD identification strategy that controls for local preference sets and partially identifies the causal effects of school assignment. Our bounds can be substantially tightened—and may even collapse to a single point—depending on the data and on how strong the first-stage assumptions are. The more restrictions we impose on strategic behavior, the more we learn about true local preferences, which then leads to narrower bounds.
@@
Strategic behavior among students depends on the characteristics of the assignment mechanism. A mechanism is said to be strategy proof if submitting true preferences is a weakly dominant strategy for all students. For example, dubins1981 demonstrate that the DA mechanism is strategy proof. However, this result breaks down when the mechanism imposes constraints on the preferences that students can submit. In many real-world school assignment mechanisms, the number of schools is too large for students to feasibly rank all schools. The central authorities running these systems may either limit the number of schools that students may rank or impose costs on the basis of the number of schools submitted (see Table 1 Panel B by fack2019 for examples).
@@
Our first-step tools for partial identification of local preferences naturally require assumptions on students’ strategic behavior. We motivate our assumptions following the important contributions of haeringer2009. They study a game where students submit constrained preference rankings and a central mechanism allocates the students to schools. One of their important findings is that it is rational for students to submit partial orders of their true preferences in some mechanisms. Specifically, suppose that a mechanism is strategy proof when students are free to rank any number of schools, as under, e.g., the unconstrained DA or Top Trading Cycles (TTC) mechanisms. Then, if the preference rankings are constrained to having at most $K$ schools, a student can do no better than selecting $K$ schools among her acceptable schools and ranking them according to her true preferences.
@@
The key assumption for our first-step tools is that students submit only partial orders of their preferences. This feature is in addition to the cutoff characterization of the matching, which we assume throughout the paper. Cutoff characterization means that a student is matched to her best feasible school, where “best” is defined according to her true preferences and a school is feasible if the student’s placement score clears that school’s admission cutoff. Our assumption about cutoff characterization is satisfied when the matching outcome is stable (azevedo2016). Stability means that each student is matched to an acceptable school and all slots in preferred schools have been filled with people who have better placement scores. In constrained DA and TTC mechanisms, stability occurs in Nash equilibrium of the preference revelation game under appropriate conditions on the placement scores (Theorems 6.3 and 6.4, haeringer2009). Stability occurs in Nash equilibrium without restrictions on the scores in the constrained serial dictatorship (SD) mechanism, which is a particular case of DA. We characterize sharp local preference sets for every individual that are compatible with the observed data and these model assumptions. For an individual drawn at random, this becomes a random set that contains the true local preference with probability one. We show how to shrink these sets by proposing a menu of behavioral assumptions on strategic behavior that researchers may find more or less appropriate to their empirical contexts. The smaller the sets are in the first step, the narrower the bounds will be in the second step, even collapsing to point identification in some cases.
@@
The second step of our approach relies on the local preference sets constructed in the first step with either our method or an alternative method. Given interest in a pair of schools $(j,k)$, we select all individuals whose local preference sets contain $(j,k)$ and whose placement scores are close to the cutoff for admission at school $j$. This subpopulation of individuals contains all individuals whose true local preferences equal $(j,k)$, but also other individuals. The average outcome in the subpopulation equals a weighted average of two averages: first, the average outcome for individuals with true local preferences $(j,k)$, which is interesting for the identification of causal effects, and second, the average outcome for individuals with true local preferences that differ from $(j,k)$. We do not know which individuals have preferences $(j,k)$, but we do characterize sharp bounds on the proportion of such individuals in the subpopulation using the random sets constructed in the first step. Thus, our setting fits the identification problem with corrupted data studied by horowitz1995. This method allows us to derive closed-form bounds on the first out of the two average outcomes above, which then leads to bounds on the average causal effects. This closed-form approach offers some intuition on when we can expect the bounds to be informative about or equal to the actual average causal effect (i.e., point identification). Although practical and intuitive, these closed-form bounds may not be sharp. Thus, building on molinari2020 and using random set theory, we characterize sharp bounds that are numerically computable when outcomes take finitely many values.
@@
Methods combining RD identification with school matching data have been popular among applied and theoretical researchers in economics for at least 15 years (kirabo_nber2009; popurquiola2011; bertanha_jmp2014). To the best of our knowledge, our paper is the first to prove RD identification of returns of school assignment in matching mechanisms with strategically reported preferences. We note that our two-step approach differs from the usual control function approach because our first step partially identifies the control variable instead of point-identifying it as in the usual approach. For this reason, we call it control mapping approach. Various other examples in econometrics feature this structure, see han_kaido_2025.
Our paper unifies and complements two branches of the literature. One branch features the methods proposed by agarwal2018 and fack2019 that take into account strategic reporting and identify students’ true preferences, but they do not focus on causal effects of school matches on future outcomes. The other branch features the methods of kirkeboen2016 and abdul_ecta2022 that identify the causal effects of different assignments but control for reported instead of true preferences. [^4] In the same branch, chen2026 provides a comprehensive set of causal parameters that are identified when assignment is based both on lottery- and RD-driven variation. An RD strategy that controls for reported preferences identifies causal parameters that could be useful for a class of counterfactuals that differs from the one we consider in this paper. It is important to note, however, that in some empirical contexts, agents may have precise ex ante information about ex post matching cutoffs at the time when they choose preferences to report. In some cases, this can even lead to a discontinuous change of submission behavior at the ex post cutoffs. We present evidence of such behavior in the empirical context of Chile (Section 5.1). Such discontinuous behavior may confound identification in an RD that controls for reported preferences; we find that controlling for true preferences makes the RD strategy robust to this possibility.
Our work, and the kind of parameters we target, open avenues for studying identification in various policy settings while allowing for strategic reporting. Recent work in causal inference seeks to identify counterfactual policy effects under interference from market equilibrium or centralized assignment (Arkhangel2025; munro_causal_2025). This line of work relies on truthful reporting but acknowledges its limitations and the importance of new developments that allow for strategic reporting.
We apply our two-step identification strategy to matching data from Chile. Chile has a centralized DA mechanism that assigns students to university–major pairs. To have an idea of the figures, in 2010, 88,000 students were constrained to rank at most eight university-major pairs ($K=8$) out of a total of 1,092 options available, in which case the theory predicts lack of truth telling. Prior work documents strategic behavior in Chile (larroucaurios2020; larroucaurios2021) and we present two additional pieces of evidence. First, ex post admission cutoffs are highly predictable ex ante for many university-major pairs (Figure 1). Second, preference submission behavior changes when students’ ex ante placement scores are close to ex post cutoffs, strongly suggesting students act on information on prior cutoffs (Figure 2). For some major pairs, the submission behavior even changes discontinuously at the ex post cutoffs (Figure 3). The methods proposed by kirkeboen2016 and abdul_ecta2022 are unable to provide the answers we seek with the Chilean data. First, strategic behavior leads their methods to identify a parameter that differs from an average treatment effect conditional on true preferences, which is our interest. Second, even if all agents report the truth, their methods are not directly applicable because of some specific aspects of the Chilean data. [^5] The methods of kirkeboen2016 apply to the SD mechanism, which is a particular case of the Chilean DA mechanism. In DA contexts, individual counterfactual sets must be defined in a more general way (see Definitions 3 and 4 below). While abdul_ecta2022 address this issue by creating a propensity-score control variable, their technique relies on placement scores with integer priorities plus continuous full-support lottery priorities. This particular structure does not correspond to the Chilean case, where program-specific placement scores are computed as functions of five primitive scores, and sometimes these functions are nonlinear. The Chilean setting requires again the counterfactual sets of schools to be carefully defined such that controlling for local preferences does not violate the continuity assumptions required by RD (see Assumption 2 and Lemma 1 below).
@@
Our analysis of the Chilean data proceeds in two parts. First, we show evidence that outcomes are heterogenous with respect to true local preferences. We compute bounds for average graduation rates if assigned to program $k$ among students with true local preference $(j,k)$. We fix the next-best alternative $k$ to be the Bachillerato de Ingreso Común at the University of Chile—a selective entry program with a large applicant pool that allows students to explore multiple disciplines before settling on a major, and is often viewed as an alternative route into competitive programs. We vary the first-best option $j$ and find non-overlapping bounds in several cases. The results reveal substantial heterogeneity across $j$, indicating that students’ preferences matter for graduation outcomes and are likely correlated with unobserved factors such as effort or ability not captured by test scores. The second part of our analysis focuses on average effects on graduation outcomes from changing the initial assignment from $k$ to $j$ among students with true local preference $(j,k)$. We choose $j$ to be Medicine at PUC Santiago because it is a highly-selective and popular program with a large applicant pool. We fix $j$ and plot bounds for common next-best alternatives $k$. In most cases, our bounds identify a positive effect of assignment into $j$ on graduating from $j$, with magnitudes that vary across alternatives $k$.
More generally, we compute bounds for three hundred college-major pairs $(j,k)$ under assumptions of varying strength on strategic behavior, which provides a valuable sensitivity exercise for researchers to evaluate the empirical content of their assumptions. We highlight three important takeaways from our bounding exercise. First, bounds are generally wide under the weakest of our assumptions, namely, that students simply select any subset with at most $K$ options from their true preference list while preserving the true relative ordering. We call this weak partial order (WPO). Wide WPO bounds essentially say that RD alone applied to matching data with strategic behavior does not reveal much unless researchers are willing to make stronger assumptions on strategic behavior. Second, bounds are generally narrow and even collapse to point identification in some cases under the strongest of our assumptions: strong partial order (SPO). SPO says that agents who rank less than $K$ programs reveal their true preferences. Third, point-estimates produced by assuming truth-telling for all agents fall outside our SPO bounds in many cases. This is relevant because the truth-telling assumption implies the SPO assumption, and we would expect point-estimates inside the bounds. Our finding suggests violation of the truth-telling assumption despite the fact that 80% of students in our sample rank less than $K$ programs. This runs counter to the common intuition that students will report truthfully whenever the $K$ constraint is non-binding for most of them.